OpenStudy (anonymous):

y=30 when x=1/6 then what is the value of k?

OpenStudy (anonymous):

whr is k but bossy??

OpenStudy (anonymous):

it does not give me a value for k

OpenStudy (anonymous):

where the equations? hide smewhere?

OpenStudy (anonymous):

Find the variation constant in which y varies inversely as x, and the following condition exists. y=30 when x=1/6 then k =?

OpenStudy (anonymous):

That is all it says

OpenStudy (anonymous):

0u.. ok.. i see it.. wait..

OpenStudy (anonymous):

the questions need the constant value..

OpenStudy (anonymous):

y=k/x 30=k/6 k=5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ohh bossy when u say y is inversely proportional then u get it